Static Mapping of Functional Programs An Example in Signal Processing

Computation Structures Group Memo 376 March 3, 1995

Jack B. Dennis

This memo is a paper presented at the conference on High Performance Functional Computing, Denver, April 1995

The research reported in this docment was performed, in part, using facilities of the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for the Laboratory is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-92-J-1310.